# Multiclass Local Calibration With the Jensen-Shannon Distance

### v1

@@ B.1.2 Proof of Theorem 3
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\|\hat{\theta}(\mathbf{\hat{p}}_{i}\mid\mathbf{x}_{i})-\mathbf{\hat{p}}_{i}\|_{1}\leq\frac{L}{|B_{b}|}\,\mathbb{E}\!\Big[\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}\Big]+\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\sqrt{\frac{2\log\!\bigl(\tfrac{n}{\delta}\bigr)}{n_{i}^{\mathrm{eff}}}}$
Conclusion: for $k\geq\max_{y\in\mathcal{Y}}y\in[1/C,1]$, averaging over bins (weight $|B_{b}|/n$) yields the final bound:
$\boxed{\Pr\Bigg(\mathrm{LCE}\leq\;k\Bigg[\varepsilon\;+\;\frac{L}{n}\sum_{b=1}^{m_{B}}\mathbb{E}\!\Big[\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}\Big]\;+\;\frac{1}{n}\sum_{b=1}^{m_{B}}\sum_{i\in I_{b}}\sqrt{\frac{2\log\!\bigl(\tfrac{n}{\delta}\bigr)}{n_{i}^{\mathrm{eff}}}}\Bigg]\Bigg)\geq 1-\delta,}$ (19)
The bound decomposes into a bias term $\mathbb{E}[Z]$, which depends on the kernel radius through the weights $w_{i,j}$, and an average variance term that scales as $1/\sqrt{n_{i}^{\mathrm{eff}}}$. Smaller kernel radii yield more concentrated weights: this reduces bias but also decreases $n_{i}^{\mathrm{eff}}$, thereby inflating the average variance. Conversely, larger kernels spread the weights more evenly, which decreases variance at the expense of bias. This captures the bias–variance tradeoff.
B.1.3 Proof of Theorem 4
@@
Then, the average Jensen-Shannon distance computed using $\hat{Q}_{i}$ converges to the one computed using the true distribution:
@@ C.1 Extension of Theorem 2 to ECCE
The subsequent analysis aims to extend the applicability of Theorem 2 to the specific class of cumulative binning-based metrics, with a focus on $ECCE$ (Expected Cumulative Calibration Error). Unlike standard binning metrics, which directly compare per-bin statistics (as defined in definition 2), cumulative binning metrics operate on the cumulative sums of per-bin statistics. Despite this systematic difference, we demonstrate that cumulative binning metrics, specifically $ECCE$, admit an upper bound of an analogous form to that presented in Theorem 2. This establishes cumulative binning metrics as a special case under the unifying bound structure provided.
Let us consider a multi-class classification setting with label space $\mathcal{Y}=\{0,\ldots,C-1\}$, and assume that the dataset $D=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ is drawn from an unknown joint distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$, with each $\mathbf{x}_{i}\in\mathbb{R}^{m}$ and $\mathbf{y}_{i}\in\{0,1\}^{C}$ being the one-hot encoding of label $y_{i}$. We consider a probabilistic classifier $f\colon\mathcal{X}\rightarrow\Delta^{C}$, where $\Delta^{C}$ is the (C1)(C-1)-dimensional probability simplex. Let $\mathbf{\hat{p}}_{i}=f(\mathbf{x}_{i})$ denote the predicted class probabilities for $\mathbf{x}_{i}$ and let $\gamma$ be the bandwidth parameter used to compute the kernel estimates $\hat{\theta}(\mathbf{y}\mid\mathbf{x})$ on a disjoint set of instances.
Let a deterministic binning function $\beta:\Delta_{c}\rightarrow\{1,\ldots,m_{B}\}$ partition the probability simplex $\Delta^{C}$ into $m_{B}$ disjoint bins $\{B_{b}\}_{b=1}^{m_{B}}$. For each bin $B_{b}$, define the index set of points that fall into it as $I_{b}=\{i:\mathbf{\hat{p}}_{i}\in B_{b}\}$. Finally, let $|B_{b}|$ denote the bin cardinality and their cumulative sums $S_{b}=\sum_{i\leq b}|B_{i}|$. The class-wise $ECCE$ is:
$\textit{class-wise }ECCE=\sum_{c=1}^{C}\pi_{c}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|.$
And, with at least probability $1-\delta\in[0,1]$, a bound of the same form of the one of Theorem 2 applies:
$\textit{class-wise }ECCE\leq\varepsilon+\sum_{b=1}^{m_{B}}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|\Psi(b,\mathcal{Y})|}},$ (22)
@@
$\sum_{c=1}^{C}\pi_{c}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|$
$=$ $\sum_{c=1}^{C}\pi_{c}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})+\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|$
$\leq\sum_{c=1}^{C}\pi_{c}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\Bigg[\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{empirical fluctuation}}+\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{miscalibration}}\Bigg].$
@@
Therefore, the summands are independent, bounded in [1,1][-1,1], and zero-mean. Hoeffding’s inequality applies to their average, yielding the desired concentration bound. Choosing $\tau_{b}=\sqrt{\tfrac{\log(2Cm_{B}/\delta)}{2|B_{i}|}}$ and applying the union bound over the $m_{B}\cdot C$ bin–class pairs yields: with probability at least $1-\delta$,
$\forall b,c:\qquad\frac{1}{|B_{i}|}\sum_{j\in I_{i}}\big(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\big)\leq\sqrt{\frac{\log(Cm_{B}/\delta)}{2|B_{i}|}}.$
@@
$\sum_{c=1}^{C}\pi_{c}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\Bigg[\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{empirical fluctuation}}+\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{miscalibration}}\Bigg]$
$\leq\sum_{c=1}^{C}\pi_{c}\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\Bigg[\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\sqrt{\frac{\log(Cm_{B}/\delta)}{2|B_{i}|}}+\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\varepsilon\Bigg]$
Since $\sum_{c}\pi_{c}=\sum_{b=1}^{m_{B}}\nicefrac{{|B_{b}|}}{{n}}=\sum_{i=1}^{b}\nicefrac{{|B_{i}|}}{{S_{b}}}=1$ , the bound simplifies as follows:
$\Pr\Bigg(\textit{class-wise }ECCE\leq\varepsilon+\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\sqrt{\frac{\log(Cm_{B}/\delta)}{2|\Psi(b,\mathcal{Y})|}}\Bigg)\geq 1-\delta$
with $\Psi(b,\mathcal{Y})=B_{i^{*}}$ and $i^{*}=\arg\min_{i\leq b}|B_{i}|$
Confidence Based Binning
@@
$\textit{class-wise }ECCE\leq\varepsilon+\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\sum_{c=1}^{C}\pi_{c}\sqrt{\frac{\log(Cm_{B}/\delta)}{2\min_{i\leq b}|B_{i,c}|}}\leq\varepsilon+\sum_{b=1}^{m_{B}}\frac{|B_{b}|}{n}\sqrt{\frac{\log(Cm_{B}/\delta)}{2\min_{c,i\leq b}|B_{i,c}|}}$ (23)
and by setting $\Psi(b,\mathcal{Y})=B_{i^{*},c^{*}}$ with $i*,c*=\arg\min_{c\in\mathcal{Y},i\leq b}|B_{i,c}|$ we obtain a bound of the same form of theorem 2. ∎
@@ C.2 Local Calibration and Proximity Bias
Proximity bias is a well-documented phenomenon in probabilistic classifiers [28], where models tend to exhibit systematic miscalibration on sparsely represented instances. This behavior is particularly concerning, as it can introduce unintended biases against underrepresented subpopulations. Addressing proximity bias is therefore critical to ensuring fairness in algorithmic decision-making, especially in high-stakes domains such as law and medicine, where equitable and reliable predictions are essential. The most effective way to characterize this phenomenon is by directly comparing the class frequency distributions of two subgroups that share similar model confidence scores but differ in input-space density.
@@
If a classifier $f$ satisfies local calibration, then with probability at least $1-\delta\in[0,1]$ the difference in class frequencies between the two sub-bins is bounded as follows:
$\Pr\left(\left\|\text{freq}(S_{1})-\text{freq}(S_{2})\right\|_{1}\leq 2\varepsilon+\sqrt{\frac{2\log(4C/\delta)}{\min(|S_{1}|,|S_{2}|)}}+\left\|\text{conf}(S_{1})-\text{conf}(S_{2})\right\|_{1}\right)\geq 1-\delta.$ (24)
A detailed proof is provided:
Proof of Theorem 5
@@
where $\mathbf{x}_{(i,1)},\ldots,\mathbf{x}_{(i,k)}$ are the $k$ nearest neighbors of $\mathbf{x}_{i}$ in feature space (excluding $\mathbf{x}_{i}$ itself). We aim to bound the quantity $\left\|\text{freq}(S_{1})-\text{freq}(S_{2})\right\|_{1}$, which measures the difference in empirical label distributions between the two sub-bins, under the assumption that the model satisfies local calibration:
$\left\|f(\mathbf{x}_{i})-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\right\|_{1}\leq\varepsilon,\quad\forall i\in\{1,\dots,n\}.$
Proof.
@@
$\Pr\left(\left\|\text{freq}(S_{s})-\text{conf}(S_{s})\right\|_{1}\leq\varepsilon+\sqrt{\frac{\log(2C/\delta^{\prime})}{2|S_{s}|}}\right)\geq 1-\delta^{\prime}.$
@@
$A$ $:=\left\{\left\|\text{freq}(S_{1})-\text{conf}(S_{1})\right\|_{1}>\varepsilon+\eta_{\delta_{1}}\right\},\quad\text{where }\eta_{\delta_{1}}:=\sqrt{\frac{\log(2C/\delta^{\prime})}{2|S_{1}|}},$
$B$ $:=\left\{\left\|\text{freq}(S_{2})-\text{conf}(S_{2})\right\|_{1}>\varepsilon+\eta_{\delta_{2}}\right\},\quad\text{where }\eta_{\delta_{2}}:=\sqrt{\frac{\log(2C/\delta^{\prime})}{2|S_{2}|}}.$
@@
$\Pr(\bar{A}\cap\bar{B})\geq 1-2\delta^{\prime}.$
Under this event, we can bound eq. 25. More precisely, both the first and third terms is bounded by $\varepsilon+\sqrt{\frac{\log(2C/\delta^{\prime})}{2n_{S_{s}}}}$, and by choosing $\delta^{\prime}=\frac{\delta}{2}$ we can conclude:
$\Pr\left(\left\|\text{freq}(S_{1})-\text{freq}(S_{2})\right\|_{1}\leq 2\varepsilon+\sqrt{\frac{2\log(4C/\delta)}{\min(|S_{1}|,|S_{2}|)}}+\left\|\text{conf}(S_{1})-\text{conf}(S_{2})\right\|_{1}\right)\geq 1-\delta.$ (26)
@@
Although limited availability of data is problematic in capturing proximity bias of a locally calibrated model, we can investigate the phenomenon from a theoretical perspective in the limit of infinite data availability. This allows us to schedule the bin width reduction, bounding the admissible score change, while keeping a sufficient bin cardinality to workaround the inherent trade-off between the two. This analysis leads to the conclusion that, in the presence of local calibration, the value of proximity bias—if it could be computed with access to infinite data—would be tightly bounded by the model’s calibration error, which is explicitly controlled by the local calibration property.
@@
Then, the bound on the proximity bias asymptotically simplifies to:
$\lim_{n\to\infty}\Pr\!\Big(\|\mathrm{freq}(S_{1})-\mathrm{freq}(S_{2})\|_{1}\leq 2\varepsilon\Big)=1.$ (27)
Thus, the empirical difference in class frequencies between proximity sub-bins becomes entirely determined by $\varepsilon$. This theoretical result allows to infer that proximity bias is directly controlled by the local calibration property of a model and in the following we provide proof:
Proof of Corollary 2.
@@
$\Pr\!\Bigg(\|\mathrm{freq}(S_{s})-\mathrm{conf}(S_{s})\|_{1}\leq\varepsilon+\sqrt{\frac{\log(2C/\delta^{\prime})}{2\min(|S_{1}|,|S_{2}|)}}\Bigg)\;\geq\;1-\delta^{\prime}$

### v2

@@ B.4 Proof of Theorem 3
$\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\big|\,\hat{\theta}({\hat{p}}_{i,c}\mid\mathbf{x}_{i})-{\hat{p}}_{i,c}\,\big|\leq\frac{L}{|B_{b}|}\,\mathbb{E}\!\Big[\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}.\Big]+\frac{1}{|B_{b}|}\sum_{i\in I_{b}}\sqrt{\frac{\log\!\bigl(\tfrac{nC}{\delta}\bigr)}{2n_{i}^{\mathrm{eff}}}}$
Conclusion: as $\sum_{c=1}^{C}1/C=1$, averaging over bins (weight $|B_{b}|/n$) yields the final bound:
$\boxed{\Pr\Bigg(\mathrm{LCE}\leq\;\varepsilon\;+\;\frac{L}{n}\sum_{b=1}^{m_{B}}\mathbb{E}\!\Big[\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}\Big]\;+\;\frac{2}{n}\sum_{b=1}^{m_{B}}\sum_{i\in I_{b}}\sqrt{\frac{\log\!\bigl(\tfrac{2nC}{\delta}\bigr)}{2n_{i}^{\mathrm{eff}}}}\Bigg)\geq 1-\delta}$ (36)
The bound decomposes into a bias term $\mathbb{E}[Z]$, which depends on the kernel radius through the weights $w_{i,j}$, and an average variance term that scales as $1/\sqrt{2n_{i}^{\mathrm{eff}}}$. Smaller kernel radii yield more concentrated weights: this reduces bias but also decreases $n_{i}^{\mathrm{eff}}$, thereby inflating the average variance. Conversely, larger kernels spread the weights more evenly, which decreases variance at the expense of bias. This captures the bias–variance tradeoff.
Class-wise metric.
In the context of class-wise $LCE$ the only difference is the introduction of a dependence between bins’ cardinalities and the classes. Then, with probability at least $1-\delta$, the bound applies with the same exact form under the following minor changes:
$\mathrm{cwLCE}\;\leq\;\varepsilon\;+\;\frac{L}{nC}\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\mathbb{E}\!\Bigg[\sum_{i\in I_{b,c}}\sum_{j\in I_{b,c}}w_{i,j}\,\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}\Bigg]\;+\;\frac{2}{nC}\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\sum_{i\in I_{b,c}}\sqrt{\frac{\log\!\bigl(\tfrac{2\,n\,C}{\delta}\bigr)}{2\,n^{\mathrm{eff}}_{i,c}}}$ (37)
B.5 Proof of Theorem 4
Statement. Let $f$ be a probabilistic classifier and let $\hat{\theta}(\cdot\mid\mathbf{x})$ be a kernel estimator of $p(\cdot\mid\mathbf{x})$ that is consistent in mean squared error. Then, under standard regularity conditions:
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\hat{\theta}(\mathbf{y}\mid\mathbf{x}_{i}\right)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\mathbf{p}_{i}\right)$ (38)
Proof.
@@ B.5 Proof of Theorem 4
Then, it suffices to show that the average Jensen-Shannon distance computed using $\hat{Q}_{i}$ converges to the one computed using the true distribution:
@@ C.1 Extension of Theorem 2 to ECCE
The subsequent analysis aims to extend the applicability of Theorem 2 to the specific class of cumulative binning-based metrics, with a focus on $ECCE$ (Expected Cumulative Calibration Error). Unlike standard binning metrics, which directly compare per-bin statistics (as defined in Definition 2), cumulative binning metrics operate on the cumulative sums of per-bin statistics. Despite this systematic difference, we demonstrate that cumulative binning metrics, specifically $ECCE$, admit an upper bound of an analogous form to that presented in Theorem 2, with slight discrepancies due to the different procedure of iteration over bins being applied. This establishes cumulative binning metrics as a special case under the unifying bound structure provided.
Let us consider a multi-class classification setting with label space $\mathcal{Y}=\{1,\ldots,C\}$, and assume that the dataset $D=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ is drawn from an unknown joint distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$, with each $\mathbf{x}_{i}\in\mathbb{R}^{m}$ and $\mathbf{y}_{i}\in\{0,1\}^{C}$ being the one-hot encoding of label $y_{i}$. We consider a probabilistic classifier $f\colon\mathcal{X}\rightarrow\Delta^{C}$, where $\Delta^{C}$ is the (C1)(C-1)-dimensional probability simplex. Let $\mathbf{\hat{p}}_{i}=f(\mathbf{x}_{i})$ denote the predicted class probabilities for $\mathbf{x}_{i}$ and let $\gamma$ be the bandwidth parameter used to compute the kernel estimates $\hat{\theta}(\mathbf{y}\mid\mathbf{x})$ on a disjoint set of instances.
Let a deterministic binning function $\beta:\Delta^{C}\rightarrow\{1,\ldots,m_{B}\}$ partition the probability simplex $\Delta^{C}$ into $m_{B}$ multidimensional (Appendix A) disjoint bins $\{B_{b}\}_{b=1}^{m_{B}}$. For each bin $B_{b}$, define the index set of points that fall into it as $I_{b}=\{i:\mathbf{\hat{p}}_{i}\in B_{b}\}$. Finally, let $|B_{b}|$ denote the bin cardinality and their cumulative sums $S_{b}=\sum_{i\leq b}|B_{i}|$. The $ECCE$ is:
$ECCE=\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\frac{S_{b}}{n}\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|$
And, with at least probability $1-\delta\in[0,1]$, a bound of similar form of the one of Theorem 2 applies:
$\Pr\Bigg(ECCE\leq\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\cdot w_{b}\Bigg(\varepsilon+\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|\Psi(b,c)|}}\Bigg)\Bigg)\geq 1-\delta$ (41)
@@
$\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\frac{S_{b}}{n}\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|$
$=$ $\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\frac{S_{b}}{n}\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})+\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|$
$\leq\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\frac{S_{b}}{n}\Bigg[\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{empirical fluctuation}}+\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{miscalibration}}\Bigg].$
@@
Therefore, the summands are independent, bounded in [1,1][-1,1], and zero-mean. Hoeffding’s inequality applies to their average, yielding the desired concentration bound. Choosing $\tau_{b}=\sqrt{\tfrac{\log(\frac{2Cm_{B}}{\delta})}{2|B_{i}|}}$ and applying the union bound over the $m_{B}\cdot C$ bin–class pairs yields: with probability at least $1-\delta$,
$\forall b,c:\qquad\frac{1}{|B_{i}|}\sum_{j\in I_{i}}\big(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\big)\leq\sqrt{\frac{\log(\frac{Cm_{B}}{\delta})}{2|B_{i}|}}.$
@@
$\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\frac{S_{b}}{n}\Bigg[\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\mathbf{1}\{y_{j}=c\}-\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{empirical fluctuation}}+\underbrace{\Bigg|\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\frac{1}{|B_{i}|}\sum_{j\in\mathcal{I}_{i}}\Bigg(\hat{\theta}_{c}(\mathbf{y}_{j}\mid\mathbf{x}_{j})-f_{c}(\mathbf{x}_{j})\Bigg)\Bigg|}_{\text{miscalibration}}\Bigg]$
$\leq\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\frac{S_{b}}{n}\Bigg[\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\sqrt{\frac{\log(\frac{Cm_{B}}{\delta})}{2|B_{i}|}}+\sum_{i=1}^{b}\frac{|B_{i}|}{S_{b}}\varepsilon\Bigg]$
Since $\sum_{i=1}^{b}\nicefrac{{|B_{i}|}}{{S_{b}}}=1$ , the bound then simplifies as follows:
$\Pr\Bigg(ECCE\leq\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\cdot w_{b}\Bigg(\varepsilon+\sqrt{\frac{\log(\frac{Cm_{B}}{\delta})}{2|\Psi(b,c)|}}\Bigg)\Bigg)\geq 1-\delta$
with $\Psi(b,c)=B_{i^{\star}}$ with $i^{\star}=\arg\min_{i\leq b}|B_{i}|$ and $w_{b}=\nicefrac{{S_{b}}}{{n}}$.
Class-wise metric.
@@
$\textit{class-wise }ECCE\leq\sum_{c=1}^{C}\sum_{b=1}^{m_{B}}\pi_{c}\frac{S_{b,c}}{n}\Bigg(\varepsilon+\sqrt{\frac{\log(\frac{Cm_{B}}{\delta})}{2\min_{i}|B_{i,c}|}}\Bigg)$ (42)
and setting $\Psi(b,c)=B_{i^{\star},c}$ with $i^{\star}=\arg\min_{i}|B_{i,c}|$ and $w_{b,c}=\nicefrac{{S_{b,c}}}{{n}}$ we obtain the bound of Theorem 2. ∎
@@ C.2 Local Calibration and Proximity Bias
Proximity bias is a well-documented phenomenon in probabilistic classifiers (Xiong et al., 2023), where models tend to exhibit systematic miscalibration on sparsely represented instances. This behavior is particularly concerning, as it can introduce unintended biases against underrepresented subpopulations. Addressing proximity bias is therefore critical to ensuring fairness in algorithmic decision-making, especially in high-stakes domains such as law and medicine, where equitable and reliable predictions are essential. The most effective way to characterize this phenomenon is by directly comparing the class frequency distributions of two subgroups that share similar model confidence scores but differ in input-space density.
@@
Let $f$ be locally calibrated with error $\varepsilon$ with respect to kernel estimates on an arbitrarily large auxiliary dataset, then conditioning on the evaluation samples $\{\mathbf{x}_{i}\}_{i=1}^{n}$ and the corresponding consistent kernel estimates, with probability at least $1-\delta\in[0,1]$ the difference in class frequencies between the two sub-bins is bounded as follows:
$\Pr\left(\left\|\text{freq}(S_{1})-\text{freq}(S_{2})\right\|_{1}\leq 2\varepsilon+\sqrt{\frac{2\log(\frac{4C}{\delta})}{\min(|S_{1}|,|S_{2}|)}}+\left\|\text{conf}(S_{1})-\text{conf}(S_{2})\right\|_{1}\right)\geq 1-\delta.$ (43)
Proof.
@@
where $\mathbf{x}_{(i,1)},\ldots,\mathbf{x}_{(i,k)}$ are the $k$ nearest neighbors of $\mathbf{x}_{i}$ in feature space (excluding $\mathbf{x}_{i}$ itself).
@@
$\Pr\left(\left\|\text{freq}(S_{s})-\text{conf}(S_{s})\right\|_{1}\leq\varepsilon+\sqrt{\frac{\log(\frac{2C}{\delta^{\prime}})}{2|S_{s}|}}\right)\geq 1-\delta^{\prime}.$
@@
$A$ $:=\left\{\left\|\text{freq}(S_{1})-\text{conf}(S_{1})\right\|_{1}>\varepsilon+\eta_{\delta_{1}}\right\},\quad\text{where }\eta_{\delta_{1}}:=\sqrt{\frac{\log(\frac{2C}{\delta^{\prime}})}{2|S_{1}|}},$
$B$ $:=\left\{\left\|\text{freq}(S_{2})-\text{conf}(S_{2})\right\|_{1}>\varepsilon+\eta_{\delta_{2}}\right\},\quad\text{where }\eta_{\delta_{2}}:=\sqrt{\frac{\log(\frac{2C}{\delta^{\prime}})}{2|S_{2}|}}.$
@@
$\Pr(\neg A\cap\neg B)\geq 1-2\delta^{\prime}.$
Under this event, we can bound Eq. 44. More precisely, both the first and third terms is bounded by $\varepsilon+\sqrt{\frac{\log(\frac{2C}{\delta^{\prime}})}{2n_{S_{s}}}}$, and by choosing $\delta^{\prime}=\frac{\delta}{2}$ we can conclude:
$\Pr\left(\left\|\text{freq}(S_{1})-\text{freq}(S_{2})\right\|_{1}\leq 2\varepsilon+\sqrt{\frac{2\log(\frac{4C}{\delta})}{\min(|S_{1}|,|S_{2}|)}}+\left\|\text{conf}(S_{1})-\text{conf}(S_{2})\right\|_{1}\right)\geq 1-\delta.$ (45)
@@
Although limited availability of data is problematic in capturing proximity bias of a locally calibrated model, we can investigate the phenomenon from a theoretical perspective in the limit of infinite data availability. This allows us to schedule the bin width reduction, bounding the admissible score change, while keeping a sufficient bin cardinality to workaround the inherent trade-off between the two. This analysis allows to conclude that, under local calibration, the value of proximity bias, if it could be computed with access to infinite data, would be tightly bounded by the model’s calibration error, which is explicitly controlled by the local calibration property.
@@
$\forall\;\xi>0:\lim_{n\to\infty}\Pr\Big(\|\mathrm{freq}(S_{1})-\mathrm{freq}(S_{2})\|_{1}\leq 2\varepsilon+\xi\Big)=1$ (46)
Thus, the empirical difference in class frequencies between proximity sub-bins is controlled by $\varepsilon$ up to vanishing stochastic fluctuations. This theoretical result allows to infer that proximity bias is directly controlled by the local calibration property of a model.
Proof.
@@
$\Pr\!\Bigg(\|\mathrm{freq}(S_{s})-\mathrm{conf}(S_{s})\|_{1}\leq\varepsilon+\sqrt{\frac{\log(\frac{2C}{\delta^{\prime}})}{2\min(|S_{1}|,|S_{2}|)}}\Bigg)\;\geq\;1-\delta^{\prime}$